OpenStudy (anonymous):

How would you solve this? If you know x +y = .5 and x(1)+y(4)=1 then what is y??

OpenStudy (mathmale):

You have various methods from which to choose to solve this system of linear equations: 1. elimination by substitution 2. elimination by addition/subtraction 3. graphing Which of this have you used before?

OpenStudy (anonymous):

Substitution.... but when I solve for x and get x= 1-y(4) and plug it into the equation don't the y(4) cancel to 0??

OpenStudy (mathmale):

By x +y = .5 and x(1)+y(4)=1 I think you mean x +y = .5 (or 2x + 2y = 1) and x+4y=1 Do you agree or disagree with my interpretations?

OpenStudy (anonymous):

No, that is correct

OpenStudy (mathmale):

Choose which variable you wish to eliminate. Then choose one of the equations and solve it for that variable.

OpenStudy (anonymous):

Wait I think I solved for y? y= 1/6? so then x= 3/2?

OpenStudy (anonymous):

Wait no that doesn't add to 1...

OpenStudy (mathmale):

you'll need to check your answer yourself, by subsitution. Again: solve ONE of the two equations for either x or y. Type your result here/

OpenStudy (mathmale):

I do agree with your y=1/6. Find x.

OpenStudy (anonymous):

thank you! I figured it out correctly. y=1/6 and x=1/3

OpenStudy (mathmale):

I'm glad for you!